# Origin of effective graviton mass from spontaneous symmetry breaking of scalar field

### v1

@@ Appendix E Degravitation
E.1 Fierz-Pauli action ($\alpha=1$)
As discussed in the previous two sections, the Fierz-Pauli action ($\alpha=1$) contains 5 DoF — two helicity-2 modes, two helicity-1 modes, and one scalar DoF. To explicitly see the degravitation mechanism, we first need to isolate the observable DoF (helicity-2 modes) from the other degrees of freedom in the action (133). After this, we integrate the extra DoF from the action.
@@ E.1 Fierz-Pauli action ($\alpha=1$)
$S=\int d^{D}x\Big{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}(h_{\mu\nu}h^{\mu\nu}-h^{2})-\frac{1}{2}m^{2}F_{\mu\nu}F^{\mu\nu}-2m^{2}(h_{\mu\nu}\partial^{\mu}V^{\nu}-h\partial_{\mu}V^{\mu})+\kappa h_{\mu\nu}T^{\mu\nu}\Big{]}.$ (166)
@@
$\begin{split}S&=\int d^{D}x\Bigg{[}\mathscr{L}_{m=0}+m^{2}\Big{[}-\frac{1}{2}h_{\mu\nu}h^{\mu\nu}+\frac{1}{4}h^{2}+V_{\mu}\Box V^{\mu}+\mathscr{N}(h-\mathscr{N})\\ &-V^{\mu}\left(\partial_{\mu}h-2\partial^{\nu}h_{\mu\nu}+2\partial_{\mu}\mathscr{N}\right)\Big{]}+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}.\end{split}$ (168)
@@
$S=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}\left(1-\frac{m^{2}}{\Box}\right)\mathscr{E}^{\mu\nu,\alpha\beta}h_{\alpha\beta}-2\mathscr{N}\frac{1}{\Box}(\partial_{\mu}\partial_{\nu}h^{\mu\nu}-\Box h)+\kappa h_{\mu\nu}T^{\mu\nu}\Big{]},$ (171)
@@
$\mathscr{E}^{\mu\nu,\alpha\beta}=\left(\frac{1}{2}(\eta^{\mu\alpha}\eta^{\nu\beta}+\eta^{\mu\beta}\eta^{\nu\alpha})-\eta^{\mu\nu}\eta^{\alpha\beta}\right)\Box-2\partial^{(\mu}\partial^{(\alpha}\eta^{\ \beta)\nu)}+\eta^{\alpha\beta}\partial^{\mu}\partial^{\nu}+\eta^{\mu\nu}\partial^{\alpha}\partial^{\beta}.$ (172)
@@
$S=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}\left(1-\frac{m^{2}}{\Box}\right)\mathscr{E}^{\mu\nu,\alpha\beta}h_{\alpha\beta}^{\prime}+2\frac{D-1}{D-2}\mathscr{N}\frac{1}{\Box-m^{2}}\mathscr{N}+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\frac{1}{\Box-m^{2}}\mathscr{N}T\Big{]}.$ (174)
@@
$S=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}\left(1-\frac{m^{2}}{\Box}\right)\mathscr{E}^{\mu\nu,\alpha\beta}h_{\alpha\beta}^{\prime}+2\frac{D-1}{D-2}\mathscr{N}^{\prime}(\Box-m^{2})\mathscr{N}^{\prime}+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\kappa\mathscr{N}^{\prime}T\Big{]}.$ (176)
@@
which shows that the presence of a mass in the theory effectively reduces the strength of the metric perturbations at low energies. The left-hand side of the above equation is the same as that of massless graviton theory (GR). This shows clearly that a massive graviton is equivalent to a filtered graviton coupled to the energy-momentum tensor $T_{\mu\nu}$ and a scalar with mass $m$ coupled with gravitational strength to the trace of energy-momentum tensor $T$. The scalar is the longitudinal mode responsible for the vDVZ discontinuity [109].
E.2 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
As the previous two sections discussed, the extended Fierz-Pauli action ($\alpha=1/2$) contains 6 DoF — two helicity-2 modes, helicity-1 modes, and scalar DoF. To explicitly see the degravitation mechanism, we first need to isolate the observable DoF (helicity-2 modes) from the other degrees of freedom in the action (133). After this, we integrate the extra DoF from the action.
@@ E.2 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
$\begin{split}S&=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}(h_{\mu\nu}h^{\mu\nu}-h^{2})-\frac{1}{2}m^{2}F_{\mu\nu}F^{\mu\nu}-2m^{2}(h_{\mu\nu}\partial^{\mu}V^{\nu}-h\partial_{\mu}V^{\mu})\\ &+\kappa h_{\mu\nu}T^{\mu\nu}-\frac{1}{4}m^{2}\Big{[}h^{2}+4h\partial_{\mu}V^{\mu}+4(\partial_{\mu}V^{\mu})^{2}\Big{]}\Bigg{]}\\ &=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)+m^{2}V_{\mu}\Box V^{\mu}-2m^{2}\left(h_{\mu\nu}\partial^{\mu}V^{\nu}-\frac{1}{2}h\partial_{\mu}V^{\mu}\right)+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}\\ &=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)+m^{2}V_{\mu}\Box V^{\mu}+2m^{2}V^{\mu}\left(\partial^{\nu}h_{\mu\nu}-\frac{1}{2}\partial_{\mu}h\right)+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}.\end{split}$ (178)
@@
$\begin{split}S&=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)-\left(\partial^{\nu}h_{\mu\nu}-\frac{1}{2}\partial_{\mu}h\right)\frac{m^{2}}{\Box}\left(\partial_{\rho}h^{\mu\rho}-\frac{1}{2}\partial^{\mu}h\right)+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}\\ &=\int d^{4}x\Bigg{[}-\frac{1}{2}\partial_{\lambda}\bar{h}_{\mu\nu}\partial^{\lambda}\bar{h}^{\mu\nu}+\partial_{\mu}\bar{h}_{\nu\lambda}\partial^{\nu}\bar{h}^{\mu\lambda}+\frac{1}{4}\partial_{\lambda}\bar{h}\partial^{\lambda}\bar{h}-\frac{1}{2}m^{2}\left(\bar{h}_{\mu\nu}\bar{h}^{\mu\nu}-\frac{1}{2}\bar{h}^{2}\right)-\partial^{\nu}\bar{h}_{\mu\nu}\frac{m^{2}}{\Box}\partial_{\rho}\bar{h}^{\mu\rho}\\ &+\kappa\bar{h}_{\mu\nu}T^{\mu\nu}-\frac{\kappa}{2}\bar{h}T\Bigg{]}\\ &=\int d^{4}x\Bigg{[}\frac{1}{2}\bar{h}_{\mu\nu}\left(1-\frac{m^{2}}{\Box}\right)\bar{\mathscr{E}}^{\mu\nu,\alpha\beta}\bar{h}_{\alpha\beta}+\kappa\bar{h}_{\mu\nu}T^{\mu\nu}-\frac{\kappa}{2}\bar{h}T\Bigg{]},\end{split}$ (181)
@@
¯$\bar{\mathscr{E}}^{\mu\nu,\alpha\beta}=\left(\frac{1}{2}(\eta^{\mu\alpha}\eta^{\nu\beta}+\eta^{\mu\beta}\eta^{\nu\alpha}-\eta^{\mu\nu}\eta^{\alpha\beta})\Box-2\partial^{(\mu}\partial^{(\alpha}\eta^{\nu)\beta)}\right).$ (182)
@@
This shows clearly that a massive graviton theory for $\alpha=1/2$ case is equivalent to a filtered graviton coupled to the energy-momentum tensor $T_{\mu\nu}$ and a scalar with mass $m$ coupled with gravitational strength to the trace of energy-momentum tensor $T$. This follows from the equation of motion of metric perturbations, which is similar to the equation (177).

### v2

@@ Appendix D Degravitation
D.1 Fierz-Pauli action ($\alpha=1$)
As discussed in the previous two sections, the Fierz-Pauli action ($\alpha=1$) contains 5 DoF — two helicity-2 modes, two helicity-1 modes, and one scalar DoF. To explicitly see the degravitation mechanism, we first need to isolate the observable DoF (helicity-2 modes) from the other DoF in the action (83). After this, we integrate the extra DoF from the action.
@@ D.1 Fierz-Pauli action ($\alpha=1$)
$\begin{split}S&=\int d^{D}x\Big{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}(h_{\mu\nu}h^{\mu\nu}-h^{2})-\frac{1}{2}m^{2}F_{\mu\nu}F^{\mu\nu}\\ &-2m^{2}(h_{\mu\nu}\partial^{\mu}V^{\nu}-h\partial_{\mu}V^{\mu})+\kappa h_{\mu\nu}T^{\mu\nu}\Big{]}.\end{split}$ (116)
@@
$\begin{split}S&=\int d^{D}x\Bigg{[}\mathscr{L}_{m=0}+m^{2}\Big{[}-\frac{1}{2}h_{\mu\nu}h^{\mu\nu}+\frac{1}{4}h^{2}+V_{\mu}\Box V^{\mu}\\ &+\mathscr{N}(h-\mathscr{N})-V^{\mu}\left(\partial_{\mu}h-2\partial^{\nu}h_{\mu\nu}+2\partial_{\mu}\mathscr{N}\right)\Big{]}+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}.\end{split}$ (118)
@@
$\begin{split}S&=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}\left(1-\frac{m^{2}}{\Box}\right)\mathscr{E}^{\mu\nu,\alpha\beta}h_{\alpha\beta}-2\mathscr{N}\frac{1}{\Box}(\partial_{\mu}\partial_{\nu}h^{\mu\nu}-\Box h)\\ &+\kappa h_{\mu\nu}T^{\mu\nu}\Big{]},\end{split}$ (121)
@@
$\begin{split}\mathscr{E}^{\mu\nu,\alpha\beta}&=\left(\frac{1}{2}(\eta^{\mu\alpha}\eta^{\nu\beta}+\eta^{\mu\beta}\eta^{\nu\alpha})-\eta^{\mu\nu}\eta^{\alpha\beta}\right)\Box\\ &-2\partial^{(\mu}\partial^{(\alpha}\eta^{\ \beta)\nu)}+\eta^{\alpha\beta}\partial^{\mu}\partial^{\nu}+\eta^{\mu\nu}\partial^{\alpha}\partial^{\beta}.\end{split}$ (122)
@@
$\begin{split}S&=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}\left(1-\frac{m^{2}}{\Box}\right)\mathscr{E}^{\mu\nu,\alpha\beta}h_{\alpha\beta}^{\prime}\\ &+2\frac{D-1}{D-2}\mathscr{N}\frac{1}{\Box-m^{2}}\mathscr{N}+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\frac{1}{\Box-m^{2}}\mathscr{N}T\Big{]}.\end{split}$ (124)
@@
$\begin{split}S&=\int d^{D}x\Big{[}\frac{1}{2}h_{\mu\nu}^{\prime}\left(1-\frac{m^{2}}{\Box}\right)\mathscr{E}^{\mu\nu,\alpha\beta}h_{\alpha\beta}^{\prime}\\ &+2\frac{D-1}{D-2}\mathscr{N}^{\prime}(\Box-m^{2})\mathscr{N}^{\prime}+\kappa h_{\mu\nu}^{\prime}T^{\mu\nu}+\frac{2}{D-2}\kappa\mathscr{N}^{\prime}T\Big{]}.\end{split}$ (126)
@@
which shows that the presence of a mass in the theory effectively reduces the strength of the metric perturbations at low energies. The left-hand side of the above equation is the same as that of massless spin-2 theory (GR). This shows clearly that a massive spin-2 modes is equivalent to a filtered graviton coupled to the energy-momentum tensor $T_{\mu\nu}$ and a scalar with mass $m$ coupled with gravitational strength to the trace of energy-momentum tensor $T$. The scalar is the longitudinal mode responsible for the vDVZ discontinuity [128].
D.2 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
As the previous two sections discussed, the extended Fierz-Pauli action ($\alpha=1/2$) contains 6 DoF — two helicity-2 modes, helicity-1 modes, and scalar DoF. To explicitly see the degravitation mechanism, we first need to isolate the observable DoF (helicity-2 modes) from the other DoF in the action (83). After this, we integrate the extra DoF from the action.
@@ D.2 Extended Fierz-Pauli action ($\alpha=1/2$) in 4-D
$\begin{split}S&=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}(h_{\mu\nu}h^{\mu\nu}-h^{2})-\frac{1}{2}m^{2}F_{\mu\nu}F^{\mu\nu}\\ &-2m^{2}(h_{\mu\nu}\partial^{\mu}V^{\nu}-h\partial_{\mu}V^{\mu})+\kappa h_{\mu\nu}T^{\mu\nu}\\ &-\frac{1}{4}m^{2}\Big{[}h^{2}+4h\partial_{\mu}V^{\mu}+4(\partial_{\mu}V^{\mu})^{2}\Big{]}\Bigg{]}\\ &=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)+m^{2}V_{\mu}\Box V^{\mu}\\ &-2m^{2}\left(h_{\mu\nu}\partial^{\mu}V^{\nu}-\frac{1}{2}h\partial_{\mu}V^{\mu}\right)+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}\\ &=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)+m^{2}V_{\mu}\Box V^{\mu}\\ &+2m^{2}V^{\mu}\left(\partial^{\nu}h_{\mu\nu}-\frac{1}{2}\partial_{\mu}h\right)+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}.\end{split}$ (128)
@@
$\begin{split}S&=\int d^{4}x\Bigg{[}\mathscr{L}_{m=0}-\frac{1}{2}m^{2}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)\\ &-\left(\partial^{\nu}h_{\mu\nu}-\frac{1}{2}\partial_{\mu}h\right)\frac{m^{2}}{\Box}\left(\partial_{\rho}h^{\mu\rho}-\frac{1}{2}\partial^{\mu}h\right)+\kappa h_{\mu\nu}T^{\mu\nu}\Bigg{]}\\ &=\int d^{4}x\Bigg{[}-\frac{1}{2}\partial_{\lambda}\bar{h}_{\mu\nu}\partial^{\lambda}\bar{h}^{\mu\nu}+\partial_{\mu}\bar{h}_{\nu\lambda}\partial^{\nu}\bar{h}^{\mu\lambda}+\frac{1}{4}\partial_{\lambda}\bar{h}\partial^{\lambda}\bar{h}\\ &-\frac{1}{2}m^{2}\left(\bar{h}_{\mu\nu}\bar{h}^{\mu\nu}-\frac{1}{2}\bar{h}^{2}\right)-\partial^{\nu}\bar{h}_{\mu\nu}\frac{m^{2}}{\Box}\partial_{\rho}\bar{h}^{\mu\rho}\\ &+\kappa\bar{h}_{\mu\nu}T^{\mu\nu}-\frac{\kappa}{2}\bar{h}T\Bigg{]}\\ &=\int d^{4}x\Bigg{[}\frac{1}{2}\bar{h}_{\mu\nu}\left(1-\frac{m^{2}}{\Box}\right)\bar{\mathscr{E}}^{\mu\nu,\alpha\beta}\bar{h}_{\alpha\beta}+\kappa\bar{h}_{\mu\nu}T^{\mu\nu}-\frac{\kappa}{2}\bar{h}T\Bigg{]},\end{split}$ (131)
@@
¯$\bar{\mathscr{E}}^{\mu\nu,\alpha\beta}=\Big{[}\frac{1}{2}(\eta^{\mu\alpha}\eta^{\nu\beta}+\eta^{\mu\beta}\eta^{\nu\alpha}-\eta^{\mu\nu}\eta^{\alpha\beta})\Box-2\partial^{(\mu}\partial^{(\alpha}\eta^{\nu)\beta)}\Big{]}.$ (132)
@@
This shows clearly that a massive spin-2 theory for $\alpha=1/2$ case is equivalent to a filtered spin-2 coupled to the energy-momentum tensor $T_{\mu\nu}$ and a scalar with mass $m$ coupled with gravitational strength to the trace of energy-momentum tensor $T$. This follows from the equation of motion of metric perturbations, which is similar to the equation (127).
Appendix E Hamiltonian analysis and counting of degrees of freedom
E.1 Linearised massless gravity
First we start with the massless case as a reference. For the sake of mathematical simplicity, we consider the metric perturbation w.r.t the Minkowski background in which the Lagrangian density can be expressed as
$\mathscr{L}_{m=0}=-\frac{1}{2}\partial_{\lambda}h_{\mu\nu}\partial^{\lambda}h^{\mu\nu}+\partial_{\mu}h_{\nu\lambda}\partial^{\lambda}h^{\mu\nu}-\partial_{\mu}h^{\mu\nu}\partial_{\nu}h+\frac{1}{2}\partial_{\lambda}h\partial^{\lambda}h,$ (133)
which can be explicitly expressed as follows
$\begin{split}\mathscr{L}_{m=0}&=\frac{1}{2}(h_{jk,0})^{2}+(h_{0k,j})^{2}-\frac{1}{2}(h_{jk,l})^{2}-h_{0k,j}h_{0j,k}-2h_{0j,k}h_{jk,0}+h_{kj,l}h_{lj,k}\\ &+h_{0k,k}h_{jj,0}-\frac{1}{2}h_{jj,0}h_{kk,0}-h_{00,l}h_{jj,l}+\frac{1}{2}h_{jj,l}h_{kk,l}+h_{0j,j}h_{kk,0}\\ &+h_{jk,j}h_{00,k}-h_{jk,j}h_{ll,k}.\end{split}$ (134)
Now we compute the conjugate momenta
$\begin{split}\Pi^{00}&=\frac{\partial\mathscr{L}}{\partial h_{00,0}}=0\\ \Pi^{0i}&=\frac{\partial\mathscr{L}}{\partial h_{0i,0}}=0\\ \Pi^{jk}&=\frac{\partial\mathscr{L}}{\partial h_{jk,0}}=h_{jk,0}-(h_{0k,j}+h_{0j,k})+\delta_{jk}(2h_{0l,l}-h_{ll,0}).\end{split}$ (135)
Then we can rearrange to see the following relation
$h_{jk,0}=\Pi_{jk}+(h_{0k,j}+h_{0j,k})-\delta_{jk}(2h_{0l,l}-h_{ll,0}).$ (136)
Taking the trace of $\Pi^{jk}$, we obtain the following relation
$\Pi_{ll}=-2h_{ll,0}+4h_{0l,l}\implies(2h_{0l,l}-h_{ll,0})=\frac{1}{2}\Pi_{ll}.$ (137)
Using the above relations, we may express the conjugate momenta as
$\begin{split}\Pi_{jk}&=h_{jk,0}-(h_{0k,j}+h_{0j,k})+\frac{1}{2}\delta_{jk}\Pi_{ll}\\ \implies h_{jk,0}&=\Pi_{jk}+(h_{0k,j}+h_{0j,k})-\frac{1}{2}\delta_{jk}\Pi_{ll}.\end{split}$ (138)
As a result, the Hamiltonian can be expressed as
$\mathscr{H}_{0}=\Pi^{\mu}h_{\mu\nu,0}-\mathscr{L}=\Pi^{jk}h_{jk,0}-\mathscr{L}.$ (139)
After substituting the previous results, we obtain the following expression for the Hamiltonian density
$\begin{split}\mathscr{H}_{0}&=\frac{1}{2}(\Pi_{jk})^{2}-\frac{1}{4}(\Pi_{ll})^{2}+2h_{0k,j}\Pi_{jk}+\frac{1}{2}(h_{jk,l})^{2}-h_{jk,l}h_{lj,k}+h_{00,l}h_{kk,l}\\ &-\frac{1}{2}h_{jj,l}h_{kk,l}-h_{jk,j}h_{00,k}+h_{jk,j}h_{ll,k}\\ &\equiv\frac{1}{2}(\Pi_{jk})^{2}-\frac{1}{4}(\Pi_{ll})^{2}+2h_{0k,j}\Pi_{jk}+\frac{1}{2}(h_{jk,l})^{2}-h_{jk,l}h_{lj,k}-h_{00}\nabla^{2}h_{kk}\\ &-\frac{1}{2}h_{jj,l}h_{kk,l}+h_{00}h_{jk,jk}+h_{jk,j}h_{ll,k},\end{split}$ (140)
where in the second line, we have done integration by parts and ignore the surface terms. Note in the above expression of the Hamiltonian density, there is no time derivative of $h_{00}$ which means it acts as a Lagrange multiplier which leads to the following constraint
$\mathscr{C}_{1}=\nabla^{2}h_{ii}-\partial_{i}\partial_{j}h_{ij}=0.$ (141)
In the similar manner, there is no time derivative of $h_{0i}$ in the expression (140) which leads to the second constraint
$\mathscr{C}_{2}=\partial_{j}\Pi_{ij}=0.$ (142)
Moreover, it can be checked quite easily that the Poisson bracket between the above two constraints vanishes $\{\mathscr{C}_{1},\mathscr{C}_{2}\}=0$. These constraints having zero Poisson bracket with all the other constraints in the theory are the first class constraints. This essentially means that the Hamiltonian in (140) is first class. Because of their symmetric nature, both $h_{ij}$ and $\Pi^{ij}$ have 6 independent components in $D=4$. Thus, together they span a $12$ dimensional phase space. Eliminating $4$ constraints implies an $8$ dimensional constraint surface. Finally, by subtracting the dimension of gauge orbits (which is again $4$, since four constraints generate four gauge invariances) we are left with a total $4$ degrees of freedom. As expected, two of them are the polarization states of the massless graviton and other two belong to their conjugate momenta.
E.2 Massive Fierz-Pauli action
Here we start with the generalized Fierz-Pauli action with the following form
$\mathscr{L}_{FP}=\mathscr{L}_{m=0}-\frac{1}{2}m^{2}(h_{\mu\nu}h^{\mu\nu}-\alpha h^{2}),$ (143)
which can further be explicitly expressed as
$\begin{split}\mathscr{L}_{FP}&=\frac{1}{2}(h_{jk,0})^{2}+(h_{0k,j})^{2}-\frac{1}{2}(h_{jk,l})^{2}-h_{0k,j}h_{0j,k}-2h_{0j,k}h_{jk,0}+h_{kj,l}h_{lj,k}\\ &+h_{0k,k}h_{jj,0}-\frac{1}{2}h_{jj,0}h_{kk,0}-h_{00,l}h_{jj,l}+\frac{1}{2}h_{jj,l}h_{kk,l}+h_{0j,j}h_{kk,0}\\ &+h_{jk,j}h_{00,k}-h_{jk,j}h_{ll,k}-\frac{1}{2}m^{2}[(1-\alpha)h_{00}^{2}-2h_{0j}^{2}+h_{jk}^{2}+2\alpha h_{00}h_{ll}-\alpha h_{ll}^{2}].\end{split}$ (144)
The conjugate momenta for the massive Lagrangian are the same as for the massless Lagrangian since the mass term does not contain any $\partial_{0}$ terms. This leads to the following Hamiltonian density
$\begin{split}\mathscr{H}&=\frac{1}{2}(\Pi_{jk})^{2}-\frac{1}{4}(\Pi_{ll})^{2}+2h_{0k,j}\Pi_{jk}+\frac{1}{2}(h_{jk,l})^{2}-h_{jk,l}h_{lj,k}+h_{00,l}h_{kk,l}\\ &-\frac{1}{2}h_{jj,l}h_{kk,l}-h_{jk,j}h_{00,k}+h_{jk,j}h_{ll,k}+\frac{1}{2}m^{2}[(1-\alpha)h_{00}^{2}-2h_{0j}^{2}+h_{jk}^{2}\\ &+2\alpha h_{00}h_{ll}-\alpha h_{ll}^{2}].\end{split}$ (145)
Note that for $\alpha=1$, $h_{0i}$ does no longer act as a Lagrange multiplier as the Hamiltonian density contains the quadratic term of $h_{0i}$, however, $h_{00}$ still acts as a Lagrange multiplier, which leads to the following constraint
$\mathscr{C}=\nabla^{2}h_{ii}-\partial_{i}\partial_{j}h_{ij}-m^{2}h_{ii}=0.$ (146)
However, this time, a secondary constraint arises from the Poisson bracket $\{H,\mathscr{C}\}$
$\mathscr{C}^{\prime}=\partial_{i}\partial_{j}\Pi_{ij}+\frac{1}{2}m^{2}\Pi_{ii}.$ (147)
This means the Hamiltonian (145) is not a first class Hamiltonian anymore. The resulting set of two constraints is second class and that removes the gauge freedom from the theory. Therefore, the theory now possesses $10$ degrees of freedom, obtained by subtracting $2$ constraints from $12$ dimensional phase space, in $D=4$. Half of them are the $5$ polarizations of the massive graviton and other $5$ belong to their conjugate momenta. On the other hand, for $\alpha=\frac{1}{2}$ (extended Fierz-Pauli action), both $h_{00},h_{0i}$ do not act as Lagrange multiplier since the Hamiltonian under that condition does also contain quadratic power of $h_{00},h_{0i}$. As a result, there are no constraint in this case, hence, the number of total degrees of freedom is $12$ out of which $6$ are the polarizations of the massive graviton and other $6$ belong to their conjugate momenta. Now out of the $6$ polarizations, one represents massive spin-2 field and other is the scalar dynamical degree of freedom, namely, $h$ the trace of the spin-2 field. Moreover, earlier, we have shown that it is not a ghost degree of freedom both using the field equations of motion and Stueckelberg trick. This shows that our previous statement regarding the number of degrees of freedom in the theory are indeed correct. As can be seen in this analysis, $\alpha=1$ reduces one extra degree of freedom compared to $\alpha=1/2$. Since the treatment is covariant and does not introduce any Lorenz-violating fields, $\alpha=1/2$ does not lead to local Lorenz breaking.